# Vizing’s Theorem in Near-Linear Time

### v1

@@ Abstract
Very recently, independently and concurrently, using randomization, this runtime bound was further improved to $\tilde{O}(n^{2})$ by [Assadi, 2024] and $\tilde{O}(mn^{1/3})$ by [Bhattacharya, Carmon, Costa, Solomon and Zhang, 2024] (and subsequently to $\tilde{O}(mn^{1/4})$ time by [Bhattacharya, Costa, Solomon and Zhang, 2024]).
We present an algorithm that computes a $(\Delta+1)$-edge coloring in $\tilde{O}(m)$ time—in fact, even $O(m\log{\Delta})$ time—with high probability, giving a near-optimal algorithm for this fundamental problem.
@@ 1 Introduction
Vizing’s original proof easily extends to an $O(mn)$ time algorithm, which was improved to $\tilde{O}(m\sqrt{n})$ in the 1980s by Arjomandi [Arj82] and Gabow et al. [GNK+85] independently. More recently, Sinnamon [Sin19] simplified the algorithms of [Arj82, GNK+85] by using randomization, while shaving off extra logarithmic factors from their runtime complexities, achieving a clean $O(m\sqrt{n})$ runtime bound. Very recently, this longstanding $O(m\sqrt{n})$ time barrier was bypassed in two recent concurrent works by Assadi [Ass24] and Bhattacharya et al. [BCC+24] which improve the runtime bound to two incomparable bounds of $\tilde{O}(n^{2})$ and $\tilde{O}(mn^{1/3})$, respectively. In a follow-up work, Bhattacharya et al. [BCSZ25] improved the $\tilde{O}(mn^{1/3})$ time bound of [BCC+24] to $\tilde{O}(mn^{1/4})$.
In this work, we resolve the time complexity of randomized $(\Delta+1)$-edge coloring (up to small logarithmic factors) by presenting a near-linear time algorithm for this problem.
@@
Remarks:
  1. 1.
    Our algorithm can be implemented to run in $O(m\log{\Delta})$ time with high probability. Nonetheless, our main focus is on polynomial runtime improvements in general, and achieving a near-linear time algorithm in particular. For this reason as well as for the sake of simplicity and transparency of our techniques, the main body of the paper presents and analyzes an algorithm that runs in $O(m\log^{3}n)$ time with high probability, which showcases our most novel ideas. A more careful implementation of the same exact algorithm, with proper adaptations, achieves $O(m\log\Delta)$ time; we defer the details of this part to Appendix A.
  2. 2.
    The $O(m\log\Delta)$ time bound that our final algorithm achieves matches the longstanding time bound for $\Delta$-edge coloring bipartite graphs [COS01] (which, to the best of our knowledge, is also the best known runtime for $(\Delta+1)$-edge coloring bipartite graphs). Our algorithm is also related to a recent line of work in [BD23, BD24, Dha24] that focuses on the $\Delta=n^{o(1)}$ case, which in particular gives a randomized $(\Delta+1)$-coloring that runs in $O(m\Delta^{4}\log\Delta)$ time with high probability.
@@ 1.1 Related Work
There has been a long line of work on fast algorithms for edge coloring that use more than $\Delta+1$ colors. It was first shown in [KS87] that an edge coloring can be computed in $\tilde{O}(m)$ time when we have $\Delta+\tilde{O}(\sqrt{\Delta})$ different colors. In addition, there are algorithms which run in linear or near-linear time for $(1+\epsilon)\Delta$-edge coloring [DHZ19, BCPS24c, EK24, BD24, Dha24] when $\epsilon\in(0,1)$ is a constant. Most recently, it was shown in [Ass24] that even a $(\Delta+O(\log n))$-edge coloring can be computed in near-linear time.
There are other studies on restricted graph classes. In bipartite graphs, a $\Delta$-edge coloring can be computed in $\tilde{O}(m)$ time [CH82, COS01]. In bounded degree graphs, one can compute a $(\Delta+1)$-edge coloring in $\tilde{O}(m\Delta)$ time [GNK+85], and it was generalized recently for bounded arboricity graphs [BCPS24b]; see also [BCPS24a, CRV24, Kow24] for further recent results on edge coloring in bounded arboricity graphs. Subfamilies of bounded arboricity graphs, including planar graphs, bounded tree-width graphs and bounded genus graphs, were studied in [CY89, CN90, CK08].
Beside the literature in the classical sequential model, considerable effort has been devoted to the study of edge coloring in various computational models in the past few years, including dynamic [BM17, BCHN18, DHZ19, Chr23, BCPS24c, Chr24], online [CPW19, BGW21, SW21, KLS+22, BSVW24, DGS24], distributed [PR01, EPS14, FGK17, GKMU18, BBKO22, CHL+20, Ber22, Chr23, Dav23], and streaming [BDH+19, SB24, CMZ24, GS24] models, among others.
@@ 1.2 Technical Overview
We now present an overview of prior approaches to $(\Delta+1)$-edge coloring and describe our techniques at a high level. For the following discussions, we will assume basic familiarity with Vizing’s proof and its underlying algorithm; see Section 2.2 for more details on Vizing’s algorithm.
Prior Approaches: A generic approach for $(\Delta+1)$-edge coloring, dating back to Vizing’s proof itself, is to extend a partial $(\Delta+1)$-edge coloring of the graph one edge at a time, possibly by recoloring some subset of edges, until the entire graph becomes colored. As expected, the main bottleneck in the runtime of this approach comes from extending the coloring to the last few uncolored edges. For instance, given a graph $G=(V,E)$ with maximum degree $\Delta$, we can apply Eulerian partition to divide $G$ into two edge-disjoint subgraphs $G=G_{1}\cup G_{2}$ with maximum degrees at most $\left\lceil\Delta/2\right\rceil$. We then find $(\left\lceil\Delta/2\right\rceil+1)$-edge colorings of the subgraphs $G_{1}$ and $G_{2}$ recursively. Directly combining the colorings of $G_{1}$ and $G_{2}$ gives a coloring of $G$ with $\Delta+3$ colors, so we have to uncolor a $2/(\Delta+3)$ fraction of the edges—amounting to $O(m/\Delta)$ edges—and try to extend the current partial $(\Delta+1)$-edge coloring to these edges. Thus, the “only” remaining part is to figure out a way to color these final $O(m/\Delta)$ edges, and let the above approach take care of the rest.
@@
As one can see, flipping long alternating paths is the major challenge in Vizing’s approach. To improve this part of the runtime, [GNK+85] designed an algorithm that groups all uncolored edges into $O(\Delta^{2})$ types depending on the two colors of the Vizing chain induced by this edge. Since all the Vizing chains of the same type are vertex-disjoint, they can be flipped simultaneously and their total length is only $O(n)$. This means that the runtime of coloring all edges of a single type can be bounded by $O(n)$ time as well. This leads to an $O(n\Delta^{2})$ time algorithm for handling all $O(\Delta^{2})$ types; a more careful analysis can bound this even with $O(m\Delta)$ time. Finally, balancing the two different bounds of $O(mn/\Delta)$ and $O(m\Delta)$ yields a runtime bound of $O(m\sqrt{n})$ for coloring $O(m/\Delta)$ edges, which leads to an $\tilde{O}(m\sqrt{n})$ time algorithm using the above framework.
There has been some very recent progress that broke through this classical barrier of $O(m\sqrt{n})$ time in [GNK+85, Arj82]. In [BCC+24], the authors proved a new useful condition for extending the coloring to uncolored edges: if the uncolored edges admit a small vertex cover, then there is also a faster algorithm for extending the coloring to them. They then show how to precondition the problem so that uncolored edges do indeed admit a small vertex cover, leading to an algorithm with $\tilde{O}(mn^{1/3})$ time. In [Ass24], the author avoided the need for Eulerian partition and recursion altogether by instead designing a new near-linear time algorithm for $(\Delta+O(\log n))$-edge coloring. This algorithm borrows insights from sublinear matching algorithms in regular bipartite graphs by [GKK10] and is thus completely different from other edge coloring algorithms mentioned above. By using this algorithm, finding a $(\Delta+1)$-edge coloring directly reduces to the color extension problem with $O((m\log n)/\Delta)$ uncolored edges (by removing the colors of a $\Theta(\log{n})/\Delta$ fraction of the edges in the $(\Delta+O(\log{n}))$-edge coloring to obtain a partial $(\Delta+1)$-edge coloring first). Applying Vizing’s procedure for these uncolored edges takes additional $\tilde{O}(mn/\Delta)=\tilde{O}(n^{2})$ time, leading to an $\tilde{O}(n^{2})$ time algorithm for $(\Delta+1)$-edge coloring. Finally, in [BCSZ25], the authors showed that a $(\Delta+1)$-coloring can be computed in $\tilde{O}(mn^{1/4})$ time, by using the algorithm of [Ass24] for initial coloring of the graph and then presenting an improved color extension subroutine with a runtime of $\tilde{O}(\Delta^{2}+\sqrt{\Delta n})$ for coloring each remaining edge; the best previous color extension time bounds were either the trivial $O(n)$ bound or the bound $\tilde{O}(\Delta^{6})$ by [Ber22].
@@
Our algorithm takes an entirely different path compared to previous work on this problem; specifically, we no longer attempt to design a faster color extension for a single edge, and instead color them in large batches, which allows for a much better amortization of runtime in coloring multiple edges. This ultimately leads to our main contribution — a new randomized algorithm for solving the aforementioned color extension problem for the last $O(m/\Delta)$ edges in $\tilde{O}(m)$ time. With this algorithm at hand, we can simply follow the Eulerian partition approach mentioned earlier and obtain an algorithm for $(\Delta+1)$-edge coloring whose runtime $T(m)$ follows the recursion $T(m)\leq 2T(m/2)+\tilde{O}(m)$; this implies that $T(m)=\tilde{O}(m)$, hence, giving us a near-linear time algorithm for $(\Delta+1)$-edge coloring. This way, we will not even need to rely on the $(\Delta+O(\log{n}))$-edge coloring algorithm of [Ass24] to color the earlier parts of the graph (although one can use that algorithm instead of Eulerian partition approach to the same effect).
@@
Color Type Reduction: Recall that the runtime of $O(n\Delta^{2})$ for the color extension algorithm of [GNK+85] is due to the fact that there are generally $O(\Delta^{2})$ types of alternating paths in the graph and that the total length of the paths of each color type is bounded by $O(n)$ edges. However, if it so happens that the existing partial coloring of the graph only involves $O(\Delta)$ color types instead, then the same algorithm will only take $O(n\Delta)=O(m)$ time (by the near-regularity assumption) and we will be done. As such, the underlying idea behind our algorithm is to modify the current partial edge coloring (without decreasing the number of uncolored edges) so that the total number of color types reduces from $O(\Delta^{2})$ to $O(\Delta)$ only.
To explore this direction, let us assume for now the input graph is bipartite, which greatly simplifies the structure of Vizing fans and Vizing chains, thus allowing us to convey the key ideas more clearly; later we highlight some of the key challenges that arise when dealing with general graphs. We shall note that it has been known since the 80s that one can $\Delta$-edge color bipartite graphs in $\tilde{O}(m)$ time [CH82, COS01]. However, the algorithms for bipartite graphs use techniques that are entirely different from Vizing fans and Vizing chains, and which do not involve solving the color extension problem at all. In particular, prior to this work, it was unclear whether one can efficiently solve the color extension problem in bipartite graphs. Therefore, the assumption of a bipartite input graph does not trivialize our goal of using Vizing fans and Vizing chains for efficiently solving the color extension problem. Additionally, we can assume that in the color extension problem, the last $O(m/\Delta)$ edges to be colored can be partitioned into $O(1)$ matchings (this guarantee follows immediately from the recursive framework we outlined earlier), and that we deal with each of these matchings separately. In other words, we can also assume that the uncolored edges are vertex-disjoint.
@@
There are two technical issues regarding this path-flipping approach. Firstly, the alternating paths $P_{u}$ and $P_{v}$ could be very long, and require a long time for being flipped. More importantly, flipping $P_{u}$ and $P_{v}$ could possibly damage other $\{\alpha,\beta\}$-type (uncolored) edges that we popularized before. More specifically, say that we have popularized a set $\Phi$ of uncolored edges. However, when popularizing the next uncolored edge $(u,v)$, it could be the case that the $\{\alpha,c_{u}\}$-alternating path $P_{u}$ is ending at a vertex $u^{\prime}$ for some edge $(u^{\prime},v^{\prime})\in\Phi$. If we flip the path $P_{u}$, then $(u^{\prime},v^{\prime})$ would no longer be of $\{\alpha,\beta\}$-type as $\alpha$ would not be missing at $u^{\prime}$ anymore.
@@
Our key observation is that when $|\Phi|$ is relatively small, most choices for an alternating path $P_{u}$ cannot be ending at edges in $\Phi$. Consider the above bad example where $P_{u}$ is a $\{c_{u},\alpha\}$-alternating path ending at $u^{\prime}$ for some $(u^{\prime},v^{\prime})\in\Phi$. Let us instead look at this from the perspective of the $\{\alpha,\beta\}$-type edge $(u^{\prime},v^{\prime})\in\Phi$. For any $(u^{\prime},v^{\prime})\in\Phi$ and any color $\gamma$, there can be at most one uncolored edge $(u,v)\notin\Phi$ whose corresponding path $P_{u}$ is the same $\{\gamma,\alpha\}$-alternating path starting at $u^{\prime}$; this is because any vertex belongs to at most one alternating path of a fixed type (here, the type $\{\gamma,\alpha\}$) and each vertex belongs to at most one uncolored edge (recall that the uncolored edges form a matching). This is also true for $\{\gamma,\beta\}$-type edges for the same exact reason. As such, ranging over all possible choices of $\gamma\in[\Delta+1]$, there are at most $O(|\Phi|\Delta)$ uncolored edges $(u,v)$ whose alternating paths could damage the set $\Phi$. Therefore, as long as the size of $\Phi$ is a $o(1/\Delta)$ fraction (or more precisely, a $O(1/\Delta)$ fraction, for a sufficiently small constant hiding in the $O$-notation) of all uncolored edges, the following property holds: a constant fraction of uncolored edges $e\notin\Phi$ can be popularized using the above method without damaging $\Phi$.
@@
This property resolves both technical issues raised earlier simultaneously. Let $\lambda$ denote the number of uncolored edges. For the second issue, as long as $|{\Phi}|=o(\lambda/\Delta)$, we can take a random uncolored edge $(u,v)\notin\Phi$ and flip $P_{u}$ and $P_{v}$ if this does not damage any already popularized edge in $\Phi$; by the observation above, this happens with constant probability. For the first issue, we can show that the expected length of alternating paths $P_{u}$ and $P_{v}$, for the random uncolored edge picked above, is $O(m/\lambda)$; indeed, this is because the number of edges colored $\alpha$ or $\beta$ is $O(n)$, hence the total length of all alternating paths with one color being fixed to $\alpha$ or $\beta$ is $O(m)$. All in all, the total time we spend to popularize a single type $\{\alpha,\beta\}$ to become a $\Theta(1/\Delta)$ fraction of all uncolored edges is $O(\lambda/\Delta\cdot m/\lambda)=O(m/\Delta)$. Since coloring edges of a single type takes $O(n)=O(m/\Delta)$ time by our earlier discussion, we can color in this way a $\Theta(1/\Delta)$ fraction of all uncolored edges in $O(m/\Delta)$ expected time. As a direct corollary, we can color all uncolored edges in $O(m\log{n})$ time, hence solving the color extension problem, in (near-regular) bipartite graphs, in near-linear time. The above argument will be detailed in the proof of Lemma 4.2 in Section 6.

### v3

@@ Abstract
Very recently, independently and concurrently, using randomization, this runtime bound was further improved to $\tilde{O}(n^{2})$ by [Assadi, 2024] and $\tilde{O}(mn^{1/3})$ by [Bhattacharya, Carmon, Costa, Solomon and Zhang, 2024] (and subsequently to $\tilde{O}(mn^{1/4})$ by [Bhattacharya, Costa, Solomon and Zhang, 2024]).
In this paper, we present a randomized algorithm that computes a $(\Delta+1)$-edge coloring in near-linear time—in fact, only $O(m\log{\Delta})$ time—with high probability, giving a near-optimal algorithm for this fundamental problem.
Contents
@@ 1 Introduction
Vizing’s original proof easily extends to an $O(mn)$ time algorithm, which was improved to $\tilde{O}(m\sqrt{n})$ in the 1980s by [Arj82] and [GNK+85] independently. More recently, the algorithms of [Arj82, GNK+85] were simplified in [Sin19], while shaving off extra logarithmic factors from their runtime complexities, achieving a clean $O(m\sqrt{n})$ runtime bound. Very recently, this longstanding $O(m\sqrt{n})$ time barrier was bypassed in two concurrent works [Ass25] and [BCC+24] which improved the runtime bound to two incomparable bounds of $\tilde{O}(n^{2})$ and $\tilde{O}(mn^{1/3})$, respectively. In a follow-up work, the $\tilde{O}(mn^{1/3})$ runtime bound of [BCC+24] to was further improved to $\tilde{O}(mn^{1/4})$ in [BCSZ25].
In this work, we resolve the time complexity of randomized $(\Delta+1)$-edge coloring up to at most a log factor by presenting a near-linear time algorithm for this problem.
@@
There is a randomized algorithm that, given any simple undirected graph $G=(V,E)$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $(\Delta+1)$-edge coloring of $G$ in $O(m\log{n})$ time with high probability.
Remarks: Several remarks on our Theorem 1.1 are in order:
  • Our algorithm in Theorem 1.1 does not rely on any of the recent developments in [Ass25, BCC+24, BCSZ25] and takes an entirely different path. We present an overview of our approach, as well as a comparison to these recent developments in Section 1.2.
  • Our main contribution in this work is to improve the time-complexity of $(\Delta+1)$-edge coloring by polynomial factors all the way to near-linear. For this reason, as well as for the sake of transparency of our techniques, we focus primarily on presenting an $O(m\log^{3}{n})$ time randomized algorithm (Theorem 6.1), which showcases our most novel ideas. Later in Section 10, we show that a more careful implementation of the same algorithm achieves a clean $O(m\log n)$ runtime.
  • We can additionally use a result of [BD23] to further replace the $\log{n}$ term in Theorem 1.1 with a $\log{\Delta}$ term, leading to an algorithm for $(\Delta+1)$-edge coloring in $O(m\log{\Delta})$ time with high probability (Corollary C.1). This matches the longstanding time bound for $\Delta$-edge coloring bipartite graphs [COS01] (which, to the best of our knowledge, is also the best known runtime for $(\Delta+1)$-edge coloring bipartite graphs). This bound is also related to a recent line of work in [BD23, BD24, Dha24b] that focused on the $\Delta=n^{o(1)}$ case and gave a randomized $(\Delta+1)$-coloring algorithm that runs in $O(m\Delta^{4}\log\Delta)$ time with high probability [BD24].
  • Vizing’s theorem generalizes for (loop-less) multigraphs, asserting that any multigraph with edge multiplicity at most $\mu$ can be $(\Delta+\mu)$-edge colored [Viz64, Viz65]. A related result is Shannon’s theorem [Sha49] that asserts that any multigraph can be $\left\lfloor 3\Delta/2\right\rfloor$ edge colored independent of $\mu$; both these bounds are tight: see the so-called Shannon multigraphs [Viz65]. We show that our techniques extend to these theorems as well, giving $O(m\log{\Delta})$ time algorithms for both problems (see Theorems A.1 and B.1 and Corollary C.1).
@@ 1.1 Related Work
In addition to algorithms for Vizing’s theorem, there has also been a long line of work on fast algorithms for edge coloring that use more than $\Delta+1$ colors. It was first shown in [KS87] that an edge coloring can be computed in $\tilde{O}(m)$ time when we have $\Delta+\tilde{O}(\sqrt{\Delta})$ different colors. In addition, there are algorithms which run in linear or near-linear time for $(1+\epsilon)\Delta$-edge coloring [DHZ19, BCPS24c, EK24, BD24, Dha24b] when $\epsilon\in(0,1)$ is a constant. Most recently, it was shown in [Ass25] that even a $(\Delta+O(\log n))$-edge coloring can be computed in $O(m\log{\Delta})$ expected time.
There are other studies on restricted graph classes. In bipartite graphs, a $\Delta$-edge coloring can be computed in $\tilde{O}(m)$ time [CH82, COS01, Alo03, GKK10]. In bounded degree graphs, one can compute a $(\Delta+1)$-edge coloring in $\tilde{O}(m\Delta)$ time [GNK+85], and it was generalized recently for bounded arboricity graphs [BCPS24b]; see also [BCPS24a, CRV24, Kow24] for further recent results on edge coloring in bounded arboricity graphs. Subfamilies of bounded arboricity graphs, including planar graphs, bounded tree-width graphs and bounded genus graphs, were studied in [CY89, CN90, CK08].
Beside the literature on classical algorithms, considerable effort has been devoted to the study of edge coloring in various computational models in the past few years, including dynamic [BM17, BCHN18, DHZ19, Chr23, BCPS24c, Chr24], online [CPW19, BGW21, SW21, KLS+22, BSVW24, BSVW25, DGS25], distributed [PR01, EPS14, FGK17, GKMU18, GP20, BBKO22, CHL+20, Ber22, Chr23, Dav23], and streaming [BDH+19, SB24, CMZ24, GS24] models, among others.
@@ 1.2 Technical Overview
We now present an overview of prior approaches to $(\Delta+1)$-edge coloring and describe our techniques at a high level. For the following discussions, we will assume basic familiarity with Vizing’s proof and its underlying algorithm; see Section 4 for more details on Vizing’s algorithm.
Prior Approaches: A generic approach for $(\Delta+1)$-edge coloring, dating back to Vizing’s proof itself, is to extend a partial $(\Delta+1)$-edge coloring of the graph one edge at a time, possibly by recoloring some edges, until the entire graph becomes colored. As expected, the main bottleneck in the runtime of this approach comes from extending the coloring to the last few uncolored edges. For instance, given a graph $G=(V,E)$ with maximum degree $\Delta$, we can apply Eulerian partitions to divide $G$ into two edge-disjoint subgraphs $G=G_{1}\cup G_{2}$ with maximum degrees at most $\left\lceil\Delta/2\right\rceil$. We then find $(\left\lceil\Delta/2\right\rceil+1)$-edge colorings of the subgraphs $G_{1}$ and $G_{2}$ recursively. Directly combining the colorings of $G_{1}$ and $G_{2}$ gives a coloring of $G$ with $\Delta+3$ colors, so we have to uncolor a $2/(\Delta+3)$ fraction of the edges—amounting to $O(m/\Delta)$ edges—and try to extend the current partial $(\Delta+1)$-edge coloring to these edges. Thus, the “only” remaining part is to figure out a way to color these final $O(m/\Delta)$ edges, and let the above approach take care of the rest.
@@
As one can see, flipping long alternating paths is the major challenge in Vizing’s approach. To improve this part of the runtime, [GNK+85] designed an algorithm that groups all uncolored edges into $O(\Delta^{2})$ types depending on the two colors of the Vizing chain induced by this edge. Since all the Vizing chains of the same type are vertex-disjoint, they can be flipped simultaneously and their total length is only $O(n)$. This means that the runtime of coloring all edges of a single type can be bounded by $O(n)$ as well. This leads to an $O(n\Delta^{2})$ time algorithm for handling all $O(\Delta^{2})$ types; a more careful analysis can bound this even with $O(m\Delta)$ time. Finally, balancing the two different bounds of $O(mn/\Delta)$ and $O(m\Delta)$ yields a runtime bound of $O(m\sqrt{n})$ for coloring $O(m/\Delta)$ edges, which leads to an $\tilde{O}(m\sqrt{n})$ time algorithm using the above framework.
There has been some very recent progress that broke through this classical barrier of $O(m\sqrt{n})$ time in [GNK+85, Arj82]. In [BCC+24], the authors speed up the extension of the coloring to uncolored edges when these edges admit a small vertex cover. They then show how to precondition the problem so that uncolored edges admit a small vertex cover, leading to a $\tilde{O}(mn^{1/3})$ time algorithm. In [Ass25], the author avoided the need for Eulerian partition and recursion altogether by instead designing a new near-linear time algorithm for $(\Delta+O(\log n))$-edge coloring. This algorithm borrows insights from sublinear matching algorithms in regular bipartite graphs by [GKK10] and is thus completely different from other edge coloring algorithms mentioned above. By using this algorithm, finding a $(\Delta+1)$-edge coloring directly reduces to the color extension problem with $O((m\log n)/\Delta)$ uncolored edges (by removing the colors of a $\Theta(\log{n})/\Delta$ fraction of the edges in the $(\Delta+O(\log{n}))$-edge coloring to obtain a partial $(\Delta+1)$-edge coloring first). Applying Vizing’s procedure for these uncolored edges takes additional $\tilde{O}(mn/\Delta)=\tilde{O}(n^{2})$ time, leading to an $\tilde{O}(n^{2})$ time algorithm for $(\Delta+1)$-edge coloring. Finally, in [BCSZ25], the authors showed that a $(\Delta+1)$-coloring can be computed in $\tilde{O}(mn^{1/4})$ time, by using the algorithm of [Ass25] for initial coloring of the graph and then presenting an improved color extension subroutine with a runtime of $\tilde{O}(\Delta^{2}+\sqrt{\Delta n})$ for coloring each remaining edge; the best previous color extension time bounds were either the trivial $O(n)$ bound or the bound $\tilde{O}(\Delta^{4})$ by [Ber22, BD24].
@@
We will no longer attempt to design a faster color extension for a single edge, and instead color them in large batches like in [GNK+85], which allows for a much better amortization of runtime in coloring multiple edges. This ultimately leads to our main technical contribution: a new randomized algorithm for solving the aforementioned color extension problem for the last $O(m/\Delta)$ edges in $\tilde{O}(m)$ time. With this algorithm at hand, we can follow the aforementioned Eulerian partition approach and obtain a $(\Delta+1)$-edge coloring algorithm whose runtime $T(m)$ follows the recursion $T(m)\leq 2T(m/2)+\tilde{O}(m)$ with high probability; this implies that $T(m)=\tilde{O}(m)$, hence, giving us a near-linear time randomized algorithm for $(\Delta+1)$-edge coloring. This way, we will not even need to rely on the $(\Delta+O(\log{n}))$-edge coloring algorithm of [Ass25] to color the earlier parts of the graph (although one can use that algorithm instead of Eulerian partition approach to the same effect).
@@
Color Type Reduction: Recall that the runtime of $O(n\Delta^{2})$ for the color extension algorithm of [GNK+85] is due to the fact that there are generally $O(\Delta^{2})$ types of alternating paths in the graph and that the total length of the paths of each color type is bounded by $O(n)$ edges. However, if it so happens that the existing partial coloring only involves $O(\Delta)$ color types instead, then the same algorithm will only take $O(n\Delta)=O(m)$ time (by the near-regularity assumption). The underlying idea behind our algorithm is to modify the current partial edge coloring (without decreasing the number of uncolored edges) so that the number of color types reduces from $O(\Delta^{2})$ to $O(\Delta)$ only.
To explore this direction, let us assume for now the input graph is bipartite, which greatly simplifies the structure of Vizing fans and Vizing chains, thus allowing us to convey the key ideas more clearly (see Section 3 for a more detailed exposition); later we highlight some of the key challenges that arise when dealing with general graphs. We shall note that it has been known since the 80s that one can $\Delta$-edge color bipartite graphs in $\tilde{O}(m)$ time [CH82, COS01]. However, the algorithms for bipartite graphs use techniques that are entirely different from Vizing fans and Vizing chains, and which do not involve solving the color extension problem at all. In particular, prior to this work, it was unclear whether one can efficiently solve the color extension problem in bipartite graphs. Therefore, the assumption of a bipartite input graph does not trivialize our goal of using Vizing fans and Vizing chains for efficiently solving the color extension problem. Additionally, we can assume that in the color extension problem, the last $O(m/\Delta)$ edges to be colored can be partitioned into $O(1)$ matchings (this guarantee follows immediately from the recursive framework we outlined earlier), and that we deal with each of these matchings separately. In other words, we can also assume that the uncolored edges are vertex-disjoint.
@@
There are two technical issues regarding this path-flipping approach. Firstly, the alternating paths $P_{u}$ and $P_{v}$ could be very long, and require a long time for being flipped. More importantly, flipping $P_{u}$ and $P_{v}$ could possibly damage other $\{\alpha,\beta\}$-type (uncolored) edges that we popularized before. More specifically, say that we have popularized a set $\Phi$ of uncolored edges. When popularizing the next uncolored edge $(u,v)$, it could be the case that the $\{\alpha,c_{u}\}$-alternating path $P_{u}$ is ending at a vertex $u^{\prime}$ for some edge $(u^{\prime},v^{\prime})\in\Phi$. If we flip the path $P_{u}$, then $(u^{\prime},v^{\prime})$ would no longer be of $\{\alpha,\beta\}$-type as $\alpha$ would not be missing at $u^{\prime}$ anymore. See Figure 1 for an illustration.
[…]
Figure 1: In this picture, we attempt to popularize edge $(u,v)$ by flipping the $\{\alpha,\gamma\}$-alternating path from $u$ and the $\{\beta,\lambda\}$-alternating path from $v$. However, flipping the $\{\alpha,\gamma\}$-alternating path from $u$ makes a previously popular edge $(u^{\prime},v^{\prime})$ unpopular as $u^{\prime}$ will not miss color $\alpha$ anymore.
@@
Our key observation is that when $|\Phi|$ is relatively small, most choices for an alternating path $P_{u}$ cannot be ending at edges in $\Phi$. Consider the above bad example where $P_{u}$ is a $\{c_{u},\alpha\}$-alternating path ending at $u^{\prime}$ for some $(u^{\prime},v^{\prime})\in\Phi$. Let us instead look at this from the perspective of the $\{\alpha,\beta\}$-type edge $(u^{\prime},v^{\prime})\in\Phi$. For any $(u^{\prime},v^{\prime})\in\Phi$ and any color $\gamma$, there can be at most one uncolored edge $(u,v)\notin\Phi$ whose corresponding path $P_{u}$ is the same $\{\gamma,\alpha\}$-alternating path starting at $u^{\prime}$; this is because any vertex belongs to at most one alternating path of a fixed type (here, the type $\{\gamma,\alpha\}$) and each vertex belongs to at most one uncolored edge (recall that the uncolored edges form a matching). This is also true for $\{\gamma,\beta\}$-type edges for the same exact reason. As such, ranging over all possible choices of $\gamma\in[\Delta+1]$, there are at most $O(|\Phi|\Delta)$ uncolored edges $(u,v)$ whose alternating paths could damage the set $\Phi$. Therefore, as long as the size of $\Phi$ is a $o(1/\Delta)$ fraction (or more precisely, an $O(1/\Delta)$ fraction, for a sufficiently small constant hiding in the $O$-notation) of all uncolored edges, the following property holds: a constant fraction of uncolored edges $e\notin\Phi$ can be popularized using the above method without damaging $\Phi$. See Figure 2 for an illustration.
@@
This property resolves both technical issues raised earlier simultaneously. Let $\lambda$ denote the number of uncolored edges. For the second issue, as long as $|{\Phi}|=o(\lambda/\Delta)$, we can take a random uncolored edge $(u,v)\notin\Phi$ and flip $P_{u}$ and $P_{v}$ if this does not damage any already popularized edge in $\Phi$; by the observation above, this happens with constant probability. For the first issue, we can show that the expected length of alternating paths $P_{u}$ and $P_{v}$, for the random uncolored edge picked above, is $O(m/\lambda)$; indeed, this is because the number of edges colored $\alpha$ or $\beta$ is $O(n)$, hence the total length of all alternating paths with one color being fixed to $\alpha$ or $\beta$ is $O(m)$. All in all, the total time we spend to popularize a single type $\{\alpha,\beta\}$ to become a $\Theta(1/\Delta)$ fraction of all uncolored edges is $O(\lambda/\Delta\cdot m/\lambda)=O(m/\Delta)$. Since coloring edges of a single type takes $O(n)=O(m/\Delta)$ time by our earlier discussion, we can color in this way a $\Theta(1/\Delta)$ fraction of all uncolored edges in $O(m/\Delta)$ expected time. As a direct corollary, we can color all uncolored edges in $O(m\log{n})$ time, hence solving the color extension problem, in (near-regular) bipartite graphs, in near-linear time. The above argument will be detailed in the proof of Lemma 3.2 in Section 3 (see also Lemma 6.3 of Section 8 for the analogous but not identical argument in general graphs).
$u^{\prime}$$v^{\prime}$$v_{1}$$u_{1}$$v_{2}$$u_{2}$$v_{3}$$u_{3}$$v_{4}$$u_{4}$$v_{5}$$u_{5}$$\beta$$\beta$$\beta$$\beta$$\beta$$\beta$$\beta$$\beta$$\beta$$\beta$
Figure 2: In this picture, $(u^{\prime},v^{\prime})\in\Phi$ with $c_{u^{\prime}}=\alpha,c_{v^{\prime}}=\beta$. For each uncolored edge $(u_{i},v_{i})$, flipping the $\{c_{i},\beta\}$-alternating path from $v_{i}$ would damage the property that $c_{v^{\prime}}=\beta$. Fortunately, there are at most $\Delta$ many different such $(u_{i},v_{i})$ because each of them is at the end of an $\{\beta,\cdot\}$-alternating path starting at $v^{\prime}$.